Question 4 0.5/2 pts You obtain a mortgage for $280,000 at a 4% interest rate with yearly payments of $16,192. Fill out the mi

i got these answers wrong! please help
Answer #1

Answer

--Mortage Schedule

Date Payment Interest Principal Loan Balance
Year 0 $280,000
Year 1 $16,192 $11,200 $4,992 $275,008

--Working

Date Payment Interest Principal Loan Balance
Year 0 280000
Year 1 16192 =280000*4% =16192-11200 =280000-4992